What would be the Km of an enzymatic reaction that showed the following kinetics?What would be the Km of an enzymatic reaction that showed the following kinetics? [substrate] mM Rxn Rate (umol/min) 8 x 10-6
Answer #1

From the graph, according to the formula,

Km= V​​​​​​max​​/2

V​​​​max= 280

Km= Vmax/2

280/2= 140

Therefore, km= 8 * 10-5 mM

Correct option A